2. Let X ~ Bin(4, ), i.e., the PMF of X is given by Px(1) = (1) (1) *,* = 0,1,2,3,4. Find Px B(k) where B = {X #0}.
Answer #1

Px]B(k) =

Where B = {X+0)

2. Let xn Bin (4 ± 2) i.e. the purf of is given PA (K) = 46: K= 0,1,2,3,4 as , Px (K) = P(X= + (x+0) i B = {x 703 = 1- P(x=0)